Should print UTKAST on the on the pages 2 and 3, light. Should print UTKAST on bottom of the page after the third page.

\usepackage[light,bottomafter]{draftcopy}
\draftcopyFirstPage{2}
\draftcopyLastPage{3}

Test7: This is a test page 1

This is a testpage 1 Test7: This is a test page 2 This is a testpage 2 Test7: This is a test page 3 This is a testpage 3 Test7: This is a test page 4 This is a testpage 4